# Negative Yields Positive: Unified Dual-Path Adapter for Vision-Language Models

### v1

@@ A.2 More Results on Domain Generalization
DualAdapter (Ours) 69.63 59.76 43.41 32.48 69.60 51.31
In Table 1 in Section 4.3 of the main text, we compare the generalization performance of our DualAdapter with other state-of-the-art methods in the presence of distribution shifts with ResNet-50 visual backbone. In Table A2, we present the performance comparison on domain generalization task using ViT-B/32 visual backbone. Consistently, our DualAdapter not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 0.57% across 4 out-of-distribution (OOD) target datasets. This verifies that our DualAdapter demonstrates superior generalizability compared to other state-of-the-art methods, independent of the visual backbone utilized.
@@ A.3 More Sensitivity Analyses of Hyper-Parameters
Building upon the sensitivity analyses of $\lambda$ and $\tau$ detailed in Section 4.4 of the main text, this section extends our examination to include the sensitivity of parameters $\alpha$ and $\beta$ on 16-shot ImageNet [4]. In our experiments on ImageNet [4], we set the hyper-parameters $\alpha$ and $\beta$ defined in Section 3 to $1.2$ and $2.0$, respectively. To comprehensively investigate the effects of different hyper-parameters, we conducted a sensitivity experiment where we varied each hyper-parameter individually and evaluated the performance on 16-shot ImageNet [4] in Fig. A3. We can see that our choice of $\alpha=1.2$ and $\beta=2.0$ yields the highest performance. Moreover, our DualAdapter maintains robust performance when adjusting these two hyper-parameters, since our DualAdapter includes adapters from both textual and visual modalities and each of our four adapters can work effectively, as we presented in Table 2 in the main text.
B Additional Implementation Details
B.1 Dataset Details
@@
In Table B4, we present the detailed statistics of each dataset we used in our experiments, including the number of classes, the sizes of training, validation and testing sets, and their original tasks.
@@
Table B4: Detailed statistics of datasets used in experiments. Note that the last 4 ImageNet variant datasets are designed for evaluation and only contain the test sets.
@@
B.2 Positive and Negative Prompts
In Table B5, we detail the specific positive and negative prompts utilized for each dataset. Additionally, as mentioned in Section 4.1, we incorporate prompts from CuPL [46] to further enhance model performance.
Table B5: Positive and negative prompts used in experiments. In addition to these prompts, we also employ CuPL [46] prompts to further enhance performance.
@@ B.2 Positive and Negative Prompts
C Future Work
In this work, we introduce the first dual-path adapter-style fine-tuning framework for VLMs. We believe that the similar concept of dual learning can also be applied to prompt-based learning methods, and be extended to fine-tune other foundational models (e.g., other VLMs [24, 74] and LLMs [63, 5]). Besides, we also notice that there are a lot of research efforts dedicated to design better prompts (e.g., using LLM [46, 40, 70]) to fully exploit the capabilities of CLIP. We hope that with our work, future research endeavors can also be directed to investigate the utilization of negative prompts to better activate the negative inference capabilities, further broadening the scope and effectiveness of CLIP.

### v2

@@ A.2 More Results on Domain Generalization
SimNL (Ours) 69.63 59.76 43.41 32.48 69.60 51.31
In Table 1 in Section 4.2 of the main text, we compare the generalization performance of our SimNL with other state-of-the-art methods in the presence of distribution shifts with ResNet-50 visual backbone. In Table A2, we present the performance comparison on domain generalization task using ViT-B/32 visual backbone. Consistently, our SimNL not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 0.57% across 4 out-of-distribution (OOD) target datasets. This verifies that our SimNL demonstrates superior generalizability compared to other state-of-the-art methods, independent of the visual backbone utilized.
@@ A.3 More Sensitivity Analyses of Hyper-Parameters
Building upon the sensitivity analyses of $\lambda$ and $\tau$ detailed in Section 4.3 of the main text, this section extends our examination to include the sensitivity of parameters $\alpha$ and $\beta$ on 16-shot ImageNet [6]. In our experiments on ImageNet [6], we set the hyper-parameters $\alpha$ and $\beta$ defined in Section 3 to $1.2$ and $2.0$, respectively. To comprehensively investigate the effects of different hyper-parameters, we conducted a sensitivity experiment where we varied each hyper-parameter individually and evaluated the performance on 16-shot ImageNet [6] in Fig. A3. We can see that our choice of $\alpha=1.2$ and $\beta=2.0$ yields the highest performance. Moreover, our SimNL maintains robust performance when adjusting these two hyper-parameters, since our SimNL includes adapters from both textual and visual modalities and each of our four adapters can work effectively, as we presented in Table 4 in the main text.
B More Discussions on Related Work
B.1 Differences between Our Proposed Method and Traditional Negative Learning
In this work, we apply the concept of negative learning to vision-language few-shot adaptation. However, our method is different with traditional negative learning approaches [20, 66]. Specifically, traditional negative learning approaches aim to learn a positive classifier using negatibe labels, while our method aim to learn a negative classifier using the positive label.
Negative Learning. In the negative learning setting [20, 66], we are given complementary labels $\bar{y}\in\mathcal{Y}\,\backslash\{y\}$, which represents the image does not belong to a specific class. Similarly, the encoded one-hot labels are given by $\bar{\boldsymbol{y}}\in\{0,1\}^{C}$. This leads to a complementary cross-entropy loss for optimizing the classifier parameters $\theta$. The training objective for $\mathcal{F}_{\theta}$ can thus be expressed as
$\min_{\theta}\mathcal{R}(\mathcal{F}_{\theta})=\mathbb{E}_{(\boldsymbol{x},\bar{y})\sim P(\boldsymbol{x},\bar{y})}\left[\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),\bar{y}\right)\right],$
$\text{where}\,\,\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),\bar{y}\right)=-\sum_{k=1}^{C}\bar{\boldsymbol{y}}_{k}\log(1-\boldsymbol{p}_{k}).$ (13)
Here, the goal is to minimize the probability corresponding to the complementary label, i.e., $\boldsymbol{p}_{\bar{y}}\rightarrow 0$.
Our Method. In this work, we apply the concept of negative learning to vision-language few-shot adaptation by employing a distinct CLIP-based negative classifier $\mathcal{G}_{\varphi}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ with parameters $\varphi$. This classifier predicts the negative probability $\bar{\boldsymbol{p}}=\texttt{Softmax}(\mathcal{G}_{\varphi}(\boldsymbol{x}))$ that the image does not belong to specific classes. The expected classification risk for $\mathcal{G}_{\varphi}$ can be written as
$\min_{\varphi}\mathcal{R}(\mathcal{G}_{\varphi})=\mathbb{E}_{(\boldsymbol{x},y)\sim P(\boldsymbol{x},y)}\left[\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)\right],$
$\text{where}\,\,\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log(1-\bar{\boldsymbol{p}}_{k}).$ (14)
By optimizing this risk, we aim to reduce the negative probability $\bar{\boldsymbol{p}}_{y}\rightarrow 0$ for the true label.
B.2 Differences between Our Proposed Method and Contrastive Learning
In this work, we propose a negative learning-based approach for vision-language few-shot adaptation. However, this “negative" refers to a negative classifier, not the negative samples pairs in contrastive learning literature. In the following discussions, we show that our negative learning is fundamentally different with contrastive learning.
Specifically, CLIP operates as a similarity-based classifier where classification logits are derived from the similarities between image features $f_{v}$ and class-specific text features $\{f_{t_{c}}\}_{c=1}^{C}$:
$\boldsymbol{p}_{y}=\frac{\exp\left(\mathrm{cos}\left(f_{t_{y}},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{v}\right)/t\right)}},$ (15)
In the context of few-shot adaptation of CLIP, both prompt-based learning methods and adapter-style fine-tuning methods seek to optimize $\{f_{t_{c}}\}_{c=1}^{C}$, whether by tuning the input prompt or directly modulating these features. As discussed in Section 3.1, these methods primarily employ positive learning, updating parameters via cross-entropy loss:
$\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),y\right)$ $=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log\boldsymbol{p}_{k}$
$=-\log\frac{\exp\left(\mathrm{cos}\left(f_{t_{y}},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{v}\right)/t\right)}},$ (16)
where $y$ is the ground-truth label for sample $\boldsymbol{x}$. This update strategy aims to optimize $p_{y}=1$. By considering $f_{t_{y}}$ and $f_{v}$ as a positive pair and $f_{t_{\neq y}}$ and $f_{v}$ as negative pairs, the positive learning process essentially becomes contrastive in nature. Specifically, Eq. (B.2) has the same form with the contrastive InfoNCE loss [34], which seeks to minimize the distance between the positive pair $f_{t_{y}}$ and $f_{v}$ while maximizing the distance between the negative pairs $f_{t_{\neq y}}$ and $f_{v}$.
However, in this work, we apply negative learning and introduce another negative classifier $\mathcal{G}_{\varphi}$. Specifically, we mine a general negative feature $f_{t_{c}}^{-}$ for each class $c$, which is absent in samples from class $c$ but present in samples from all other classes. Then the probability that the image not belongs to class $y$ can be written as
$\bar{\boldsymbol{p}}_{y}=\frac{\exp\left(\mathrm{cos}\left(f_{t_{y}}^{-},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}}^{-},f_{v}\right)/t\right)}}$ (17)
As discussed in Section 3.1, we update the set of negative features by
$\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)$ $=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log(1-\bar{\boldsymbol{p}}_{k})$
$=-\log\left(1-\frac{\exp\left(\mathrm{cos}\left(f_{t_{y}}^{-},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}}^{-},f_{v}\right)/t\right)}}\right).$ (18)
In the language of contrastive learning, we consider all $f_{t_{\neq y}}^{-}$ and $f_{v}$ as positive pairs and $f_{t_{y}}^{-}$ and $f_{v}$ as the only negative pair.
In summary, the concept of “negative" is different in our negative learning and contrastive learning: (1) In our negative learning, “negative" specifically refers to a negative classifier. We explicitly train another negative classifier $\mathcal{G}_{\varphi}$ to ensemble with the positive classifier; (2) In contrastive learning, “negative" refers to the negative sample pairs that are constructed and utilized during training. Specifically, by constructing positive and negative sample pairs, the objective of each classifier can be formulated as a contrastive objective as discussed above.
B.3 Differences with Other Similar Approaches
To the best of our knowledge, the following works adopt similar negative learning methods to open-set problems:
  • RPL [3] is the seminal work that learns a set of reciprocal points as negative representations of target classes to enhance model’s recognition capabilities when handling unseen samples.
  • CLIPN [56] leverages additional large-scale datasets to a fine-tune a “no" text encoder to enhance CLIP’s out-of-distribution (OOD) detection capability.
  • LSN [32] also focus on learning some negative prompts for OOD detection task using CLIP.
However, we formulate negative learning from a different and more fundamental perspective, i.e., using negative prediction to directly improve the accuracy of positive prediction in a closed-set problem. We also empirically validate that leveraging negative cues from CLIP can effectively improve both few-shot classification performance and generalization capability.
C Additional Technical Details
C.1 Dataset Details
@@
In Table C4, we present the detailed statistics of each dataset we used in our experiments, including the number of classes, the sizes of training, validation and testing sets, and their original tasks.
@@
Table C4: Detailed statistics of datasets used in experiments. Note that the last 4 ImageNet variant datasets are designed for evaluation and only contain the test sets.
@@
C.2 Positive and Negative Prompts
In Table C5, we detail the specific positive and negative prompts utilized for each dataset. Additionally, as mentioned in Section 4.1, we incorporate prompts from CuPL [36] to further enhance model performance.
Table C5: Positive and negative prompts used in experiments. In addition to these prompts, we also employ CuPL [36] prompts to further enhance performance.
@@ C.2 Positive and Negative Prompts
D License Information
Datasets. We list the known license information for the datasets below:
  • MIT License: ImageNet-A [16], ImageNet-V2 [40], ImageNet-R [15], and ImageNet-Sketch [55].
  • CC BY-SA 4.0 License: OxfordPets [35].
  • Research purposes only: ImageNet [6], StandfordCars [24], DTD [5], FGVCAircraft [30], SUN397 [63].
Code. In this work, we also use some code implementations from existing codebase: CLIP [38], CoOp [74], APE [76], and CuPL [36]. The code used in this paper are all under the MIT License.
E Further Discussions
Future Work. In this work, we introduce negative learning for adapting VLMs to downstream tasks. We believe that the similar concept of negative learning can also be applied to prompt-based learning methods, and be extended to fine-tune other foundational models (e.g., other VLMs [21, 64] and LLMs [51, 7]). Besides, we also notice that there are a lot of research efforts dedicated to design better prompts (e.g., using LLM [36, 31, 62]) to fully exploit the capabilities of CLIP. We hope that with our work, future research endeavors can also be directed to investigate the utilization of negative prompts to better activate the negative inference capabilities, further broadening the scope and effectiveness of CLIP.
Limitations. We identify two potential limitations of our SimNL: (1) Its efficacy in zero-shot situations is constrained due to the scarcity of original negative descriptions in the CLIP training corpus; (2) Similar to other adapter-style fine-tuning approaches, our SimNL fine-tuned on a specific task cannot be directly applied to another task without additional adaptation. However, recently, Wang et al. [59] shows that adapter-style fine-tuning methods can be extended for these scenarios using the $k$NN algorithm.
Broader Impacts. In this work, we aim to build more reliable machine learning systems by leveraging the extensive knowledge of current foundational models. Specifically, we introduce negative learning to more efficiently transfer pre-trained VLMs to downstream tasks, enhancing both the task-specific performance of CLIP and its robustness to natural distribution shifts. Additionally, we explore the few-shot adaptation of VLMs in a noisy setting, which better aligns with real-world scenarios where mislabeled samples may exist in the support set. We hope this work inspires future studies to focus on the generalization and robustness of pre-trained large-scale foundation models.